# LaserSAM: Zero-Shot Change Detection Using Visual Segmentation of Spinning LiDAR

### v1

@@ Abstract
This paper presents an approach for applying camera perception techniques to spinning LiDAR data. To improve the robustness of long-term change detection from a 3D LiDAR, range and intensity information are rendered into virtual perspectives using a pinhole camera model. Hue-saturation-value image encoding is used to colourize the images by range and near-IR intensity. The LiDAR’s active scene illumination makes it invariant to ambient brightness, which enables night-to-day change detection without additional processing. Using the colourized, perspective range image allows existing foundation models to detect semantic regions. Specifically, the Segment Anything Model detects semantically similar regions in both a previously acquired map and live view from a path-repeating robot. By comparing the masks in both views, changes in the live scan are detected. Results indicate that the Segment Anything Model is capable of accurately capturing the shape of arbitrary changes introduced into scenes. The system achieves an object recall of 82.6% and a precision of 47.0%. Changes can be detected through day-to-night illumination variations reliably. After pixel-level masks are generated, the one-to-one correspondence with 3D points means that the 2D masks can be directly used to recover the 3D location of the changes. Eventually, the detected 3D changes can be avoided by treating them as obstacles in a local motion planner.
@@
Figure 1: The Clearpath Warthog robot driving at night during snowfall, repeating a path that was previously taught during daytime. This paper proposes LaserSAM to detect and mask environmental changes between teach and repeat paths by creating virtual camera views from LiDAR data and applying modern deep-learning-based segmentations.
@@ I Introduction
Robust perception remains a central challenge for mobile robots operating in unstructured environments. Unlike autonomous cars, there are no rules of the road and rare classes of obstructions occur with a higher frequency [1]. Occlusion effects caused by vegetation degrade the quality of many detectors trained in structured environments [2]. These factors make it difficult for supervised perception pipelines to accurately identify hazards reliably enough to maintain the aggressive performance of the vehicles. In addition to obstacle detection, off-road vehicles must perform terrain assessment to understand which local paths are feasible to drive [3]. Visual Teach and Repeat (VT&R) [4] provides a practical alternative to terrain assessment. A human operator manually drives a robot along a network of connected paths. While driving, the robot constructs local submaps, which are used later for localization. After the paths are taught, the robot can autonomously navigate between any two locations in the network. The primary responsibility of terrain assessment is delegated to the human pilot, allowing aggressive maneuvers to be demonstrated to the robot. To allow safe repeats over long periods of time, we propose that it is sufficient to detect changes in the scene that intersect the planned local path of the vehicle [5], [6]. During the teach pass, the operator defines a safe corridor [7] around the path. When repeating, the robot navigates around new obstacles within the corridor’s boundary. By focusing on detecting changes instead of specific types of hazards, any type of obstacle can be detected and avoided without predefining specific classes. For this reason, a robust change-detection module is desired to improve the system’s capabilities.
@@
Spinning $360\degree$ LiDAR scanners have emerged as a popular, complementary sensor to cameras [14]. LiDAR provides accurate 3D position and intensity measurements for each point. Most automotive LiDAR units operate at a wavelength of 1550 nm because the atmosphere absorbs most of the sun’s energy at that wavelength [15]. Ouster LiDARs operate at 840 nm, which measures significant amounts of ambient light [15]. The combination of ambient and projected infrared light leads to more natural illumination effects that blur the line between passive camera and active LiDAR.
This work blends the strengths of both modalities by rendering perspective camera images from LiDAR scans. This approach allows computer vision algorithms to be applied to multi-view change detection. Specifically, the Segment Anything Model (SAM) [16] and the Fast Segment Anything Model (FastSAM) [17] are used to detect semantic regions in the rendered images. Multi-modal algorithms rely on extrinsic calibration between LiDAR and cameras to define the depth of some pixels in a camera image. In this approach, every pixel in the generated image corresponds exactly to a 3D point. 3D points from the map and the live scan are rendered into a common virtual camera frame for analysis. Finally, the active illumination and signal processing of the 840 nm IR means that change detection works across illumination conditions with no additional processing required. Figure 1 demonstrates the types of conditions that challenge camera-based detection but do not impact the proposed LiDAR pipeline. In summary, we propose the following contribution: a change detection pipeline that leverages the sensor benefits of LiDAR and pre-trained foundation models for image segmentation.
@@ II-A Change Detection
Detecting changes in images, point clouds, or other rich sensors is often a key task in scene understanding [8]. One of the central challenges is to accurately capture the domain-specific semantics. Differentiating inconsequential changes such as illumination and sensor orientation allows for meaningful downstream processing. Ultimately, these definitions must be determined at the problem level but common binary definitions of changed and unchanged are common in the literature [8]. Most methods operate at the point/pixel level. Aggregation into objects may not be necessary or occur as a downstream processing task.
@@ II-B Zero-Shot Semantic Segmentation
In 2023, Meta Research released the Segment Anything Model (SAM) [16], which aims to perform semantic segmentation on any type of scene. The model is trained on over eleven million images with more than one billion masks. This vast dataset has enabled the capability to segment regions in a class-agnostic manner. In contrast to common datasets for autonomous driving such as CityScapes [25] or SemanticKITTI [26], this allows the model to generate masks on tasks and cameras on which it was not trained. Additionally, the model decoder can accept additional prompts such as points of interest or bounding boxes that refine the masks that are reported. While this performance is impressive, its run time is currently slow for robotics applications.
In response to this runtime constraint, FastSAM [17] was developed based on the YoloV8 [27] architecture. It is trained on the same dataset as Meta’s SAM but the simpler convolutional neural network architecture is much faster than the transformer model used originally. However, the key tradeoff for the speed boost is prompt quality. FastSAM provides a list of masks and confidences but does not consider any additional prompts such as position or bounding boxes. The impacts of this trade-off will be discussed later in the paper.
@@ II-C LiDAR as a Camera
The most common LiDAR sensors used in modern autonomous vehicle development are 360°$\degree$ units from Ouster (Velodyne) [28], and RoboSense [29]. These sensors are characterized by higher resolution along the horizontal spinning axis than in the vertical field of view. The current state-of-the-art sensors have 128 vertical beams and operate at 10 to 20 Hz [28]. However, low-framerate LiDAR sensors have existed with much higher resolutions for more than two decades. McManus et al. [30] showed that lighting-invariant visual odometry could be performed using SURF features on intensity images taken from an Autonosys 2D scanning LiDAR. This sensor produces images with a $30\degree\text{V}\times 90\degree$H FOV, with resolution $480\times 360$ pixels at 2 Hz. The framerate is much slower, but the vertical resolution is still significantly higher (12 pixels per degree) than the OS-1 LiDAR (2.85 pixels per degree) used in this paper. This approach was extended to account for the motion distortion caused by the moving sensor [31] and used to control an offroad vehicle in closed-loop [32], [33].
@@ III-A Creating Virtual Camera Views
The core idea of this pipeline is to project the two aligned point clouds into images, the format expected by SAM. The natural data format of a spinning LiDAR is equirectangular where each pixel corresponds to a constant angular offset. The natural data format of cameras is the perspective projection. Sample images created using both projections are shown in Figure 3. A larger blindspot around the base of the robot is created in the equirectangular view than in the perspective view when deviating from the path (the map view). In the following, we will use the perspective projection because the image segmentation neural networks were trained on data in this form. We define a virtual camera pose in the vehicle’s frame, $\textbf{T}_{cv}\in SE(3)$, which is used to render the images. The desired virtual camera pose is arbitrary, but aligning its position with the origin of the LiDAR sensor minimizes the amount of interpolation required in the final image. The first step is to align the point clouds in the camera frame. When a new live scan is received, an existing ICP-based localization module [34] is used to extract the relative pose of the vehicle in the map ($\textbf{T}_{vm}\in SE(3)$). The extrinsic transformation between the LiDAR sensor and the vehicle, $\textbf{T}_{vs}\in SE(3)$, is known from calibration.
@@
When a perspective projection is used, gaps in the image are created due to the distortion. These gaps are highlighted in the third row of Figure 3. Interpolating these regions is critical for the segmentation to work properly. A $3\times 3$ kernel is used to interpolate the missing pixels, but only non-zero pixels are considered. The distortion is exacerbated by offsetting the virtual camera’s origin from the LiDAR. Conceptually, the virtual camera could be colocated with the LiDAR’s position in either the teach or repeat. However, we choose to locate it in the repeat frame because the map point clouds are an accumulation of points from sequential scans. The hue-saturation-value colour model is used to colour the images by range. The $z$ coordinate in the camera frame is used for hue, mapping over a maximum range of 50 m. The saturation is the constant 255 for all pixels and the value is the LiDAR intensity of the pixel. The first column of Figure 4 shows two sample images rendered from the robot in a common frame.
@@ III-B Segmenting Regions of Interest
After rendering both images in a common frame, segmentation occurs. We compare the performance of SAM and FastSAM for this step. SAM can perform segmentation based on point prompts of interest. To generate point prompts, the norm of the pixel-wise difference of the two images is evaluated. The top $k$ local maxima of the resulting 2D surface are used as prompts to the model. The model is prompted twice, once using the live image and once using the map image. This generates two lists, each containing $k$ masks from the live and map images, respectively.
FastSAM cannot accept prompts in the same way as SAM. Instead, the model is evaluated independently for each image. This means that there are two lists of masks at each timestep. The two lists of masks are independent meaning their lengths often differ and a correspondence problem presents itself to associate masks in teach and repeat.
@@ III-C Detecting Changes
After independently segmenting the images, the detected masks are compared. To decide if two masks capture the same object, the intersection-over-union (IoU) score is calculated between them. The bitwise AND ($\wedge$) and OR ($\vee$) operations are used to calculate the IoU of two binary masks ($b1,b2$),
@@
An IoU score of one indicates a perfect match and zero means no overlap between the individual masks. Masks with a maximum IoU of less than 0.5 are considered to be changed between the live scan and the map. The remaining masks are considered unchanged. For SAM, this is a direct comparison of the matched lists. For FastSAM, a brute-force comparison between every mask in the two scans is required. There are typically fewer than ten masks so this is computationally acceptable. All points classified as changed in the live scan are then back-projected into 3D. The 3D bounding box, centroid, or raw points can be estimated and provided to a local planner. A corridor-constrained sampling-based planner [7] inflates the changed points to account for the robot’s footprint and avoids them.
TABLE I: Comparison of SAM, FastSAM, and Pixel Difference Change Detection Algorithms.
Method Pixel-Wise IoU Object Precision Object Recall Object F-Score Run Time (ms)
Baseline Pixel Difference 20.7% 10.4% 53.0% 17.4% 5.7 $\pm$ 1.2
[…]
SAM (sam_vit_b) 25.9% 47.0% 82.6% 59.9% 322 $\pm$ 35
@@ IV Experiments
Figure 4: Change-detection masks for a sample frame. The left column shows the input images. There are three changed objects, marked in the ground truth. The segmentation results for each algorithm are shown in their respective panel.
Experiments were performed at the University of Toronto’s Institute for Aerospace Studies using a Clearpath Warthog UGV [36] equipped with an Ouster OS-1 128 beam LiDAR [28]. Data were recorded during night and day as well as during snowstorm conditions. Static obstacles placed on and around the original path and pedestrians walking near the robot were introduced as changes in the dataset. A mixture of natural and fabricated items was used to minimize the impact of the chosen materials’ reflectivity on the experiments. Additionally, duplicate items were placed near the path as part of the static scene. This demonstrates that the algorithm is performing change detection, not obstacle classification. Two different sequences were recorded with changes. The first sequence is 262 m long on relatively flat terrain with cones, mannequins, and road signs as obstacles. The second sequence is 230 m long through wooded areas with pedestrians and a single cone obstacle. The forest sequence is notable because the teach pass occurred during daylight and the repeat pass was recorded in the dark, about six hours later.
The dataset has 7386 live scans to use for testing. To provide an initial, quantitative evaluation of the change detection quality, 92 frames were manually annotated to provide ground truth masks. The frames were randomly selected, using a sampling bias towards frames containing changes. The ground-truth masks were instance segmented to allow for per-object metrics to be evaluated in addition to per-pixel ones. As discussed in the related work, the precise semantic meaning of a change is problem-dependent. In this dataset, only new or moved obstacles that would impact the robot’s ability to drive were marked as changed. Small scene changes such as footprints in the snow, or tracks from previous drives were marked as unchanged. Additionally, we did not segment the teach images to mark areas where an object was removed.
@@ IV-A Results
Figure 5: A false positive detection due to misalignment. Localization has an angular error that leads to an inaccurately large pixel difference. The prompt point (red star) exists in a region that crosses the object boundary between the live and map images.
The pipeline was evaluated on the annotated subset of the frames. Specifically, SAM and FastSAM were compared to the baseline pixel-wise difference algorithm, which was used to create the point prompts for SAM. Two metrics are used for comparison: the pixel-wise IoU between the predicted and ground-truth segmentation masks and the F-score of objects being detected. The ground truth is instance-segmented, which allows for the evaluation to consider individual changes in the same frame, and whether or not they are detected. An overlap of 50% IoU is used as the threshold for a true positive detection for the object-based metrics. After counting the number of true positives (TP), false positives (FP), and false negatives (FN) detections, the precision, recall, and F-score can be evaluated [37]. The achieved results are summarized in Table I. All computational timing was performed on a laptop NVIDIA RTX A4500 GPU.
Figure 6: A temporal sequence of the semantic regions generated by FastSAM [17]. The mask colours are randomly generated. The green arrows highlight consistent mask regions through time. Red arrows point to masks that flicker in time, which is a challenge for stable processing.
Figure 4 shows the segmentation result for a sample frame with a moving pedestrian, static cone and static mannequin. All three methods correctly find the three objects but FastSAM marks extra objects, and SAM has the most accurate segmentation boundary. In our experiments, SAM outperforms both other methods by a significant margin.
@@
Figure 7: A temporal sequence of semantic regions generated by SAM [16]. The shared prompt point is highlighted in the pink bulls-eye. The mask colours match between the teach and repeat. Prompts often exist near the boundaries, which is undesirable.
In its current state, the system has a high recall, but low precision; there are too many false positives. Qualitative analysis provides some additional insight into the results. SAM and FastSAM tend to over-segment in failure cases. The most commonly observed failure mode during our experiments is a localization error, causing a large difference on the static object region boundary. In Figure 5, live and map images are not prompted inside the same object so SAM segments large, non-overlapping regions. This has an outsized impact on the IoU because many false-positive pixels are added to the IoU calculation. On the object level, the results are more encouraging. In successful frames, such as Figure 7, the changes are captured accurately and there are no isolated bright pixels common in the baseline or 3D point clouds approaches. We anticipate that with improved localization, the number of large, static false-positive detections will fall rapidly.
Figure 6 provides a sample set of three matched teach and repeat frames extracted using FastSAM. Critically, the changed objects are classified correctly within the scene. However, some masks flicker and are not retained as desired. Sometimes the ground is not masked as a region in the teach pass leading to an error where the entire ground may be considered a change in the repeat. While post-processing filters this out based on size, ground errors near the robot have the largest impact on the path planner downstream.
@@
Figure 7 shows an example sequence of masks generated by SAM. When changes are detected, the map image often has a large mask corresponding to the ground or background. Once the two masks are generated, the same IoU threshold comparison is used, as with FastSAM. The masks from SAM tend to be more consistent over time. This is advantageous for robot motion planning.
The ambient illumination of the scene does not affect detection quality. Figure 8 shows the two virtual images in the afternoon and night. The colouring of the images is qualitatively indistinguishable. This is a significant benefit compared to using a conventional camera with the same approach. Qualitatively, it appears that SAM finds accurate segmentation boundaries, which can be challenging near interfaces.

### v2

@@ Abstract
This paper presents an approach for applying camera perception techniques to spinning LiDAR data. To improve the robustness of long-term change detection from a 3D LiDAR, range and intensity information are rendered into virtual perspectives using a pinhole camera model. Hue-saturation-value image encoding is used to colourize the images by range and near-IR intensity. The LiDAR’s active scene illumination makes it invariant to ambient brightness, which enables night-to-day change detection without additional processing. Using the range-colourized, perspective image allows existing foundation models to detect semantic regions. Specifically, the Segment Anything Model detects semantically similar regions in both a previously acquired map and live view from a path-repeating robot. By comparing the masks in both views, changes in the live scan are detected. Results indicate that the Segment Anything Model accurately captures the shape of arbitrary changes introduced into scenes. The proposed method achieves a segmentation intersection over union of 73.3% when evaluated in unstructured environments and 80.4% when evaluated within the planning corridor. Changes can be detected reliably through day-to-night illumination variations. After pixel-level masks are generated, the one-to-one correspondence with 3D points means that the 2D masks can be used directly to recover the 3D location of the changes. The detected 3D changes are avoided in a closed loop by treating them as obstacles in a local motion planner. Experiments on an unmanned ground vehicle demonstrate the performance of the method.
@@
Figure 1: The Clearpath Warthog robot driving at night during snowfall, repeating a path that was previously taught during daytime. This paper proposes LaserSAM to detect and mask environmental changes between teach and repeat paths by creating virtual camera views from LiDAR data and applying deep-learning-based segmentations.
@@ I Introduction
Robust perception remains a central challenge for mobile robots operating in unstructured environments. Unlike autonomous cars, there are no rules of the road and out-of-distribution classes of obstructions occur with a higher frequency [1]. Occlusion effects caused by vegetation degrade the quality of many detectors trained in structured environments [2]. These factors make it difficult for supervised perception pipelines to identify hazards reliably enough to maintain the aggressive performance of the vehicles. In addition to obstacle detection, off-road vehicles must perform terrain assessment to understand which local paths are feasible to drive [3]. Visual Teach and Repeat (VT&R) [4] provides a practical alternative to terrain assessment. A human operator manually drives a robot along a network of connected paths. While driving, the robot constructs local submaps, which are used later for localization. After the paths are taught, the robot can autonomously navigate between any two locations in the network. The primary responsibility of terrain assessment is delegated to the human pilot, allowing aggressive maneuvers to be demonstrated to the robot. To allow safe repeats over long periods of time, we propose that it is sufficient to detect changes in the scene that intersect the planned local path of the vehicle [5], [6]. During the teach pass, the operator defines a safe corridor [7] around the path. When repeating, the robot navigates around new obstacles within the corridor’s boundary. By focusing on detecting changes instead of specific types of hazards, any type of obstacle can be detected and avoided without predefining specific classes. For this reason, a robust change-detection module is desired to improve the system’s capabilities.
@@
Spinning $360\degree$ LiDAR scanners have emerged as a popular, complementary sensor to cameras [14]. LiDAR provides accurate 3D position and intensity measurements for each point. Most automotive LiDAR units operate at a wavelength of 1550 nm because the atmosphere absorbs most of the sun’s energy at that wavelength [15]. Ouster LiDARs operate at 840 nm allowing them to capture ambient sunlight [15]. The combination of ambient and projected infrared light leads to more natural illumination effects that blur the line between passive camera and active LiDAR.
This work blends the strengths of both modalities by rendering perspective camera images from LiDAR scans. This approach allows computer vision algorithms to be applied to multi-view change detection. Specifically, the Segment Anything Model (SAM) [16] is used to detect semantic regions in the rendered images. Its zero-shot generalization capabilities allow for the segmentation of any object, including unseen ones in new domains. Multi-modal algorithms rely on extrinsic calibration between LiDAR and cameras to define the depth of some pixels in a camera image. In this approach, every pixel in the generated image corresponds exactly to a 3D point. 3D points from the map and the live scan are rendered into a common virtual camera frame for analysis. Finally, the active illumination and signal processing of the 840 nm IR means that change detection works across illumination conditions with no additional processing required. Figure 1 demonstrates the types of conditions that challenge camera-based detection but do not impact the proposed LiDAR pipeline.
In summary, we propose the following contribution: a change detection pipeline that combines the sensor benefits of LiDAR with a pre-trained foundation model for image segmentation.
@@ II-A Change Detection
Detecting changes in images, point clouds, or other rich sensors is often a key task in scene understanding [8]. One of the central challenges is to capture the domain-specific semantics accurately. Differentiating inconsequential changes such as illumination and sensor orientation allows for meaningful downstream processing. Ultimately, these definitions must be determined at the problem level but the binary definitions of changed and unchanged are common in the literature [8]. Most methods operate at the point/pixel level. Aggregation into objects may not be necessary or occur as a downstream processing task.
@@ II-B Zero-Shot Semantic Segmentation
In 2023, Meta Research released the Segment Anything Model (SAM) [16], which aims to perform semantic segmentation on any scene. The model is trained on over eleven million images with more than one billion masks. This vast dataset has enabled the capability to segment regions in a class-agnostic manner. In contrast to common datasets for autonomous driving such as CityScapes [23] or SemanticKITTI [24], this allows the model to generate masks for images that were captured in new environments with different cameras. Additionally, the model decoder accepts geometric prompts such as points of interest or bounding boxes that refine the masks that are reported. While this performance is impressive, its run time is slow for robotics applications.
@@ II-C LiDAR as a Camera
The most common LiDAR sensors used in autonomous vehicle development are 360°$\degree$ units from Ouster (Velodyne) [25], and Hesai [26]. These sensors are characterized by higher resolution along the horizontal spinning axis than in the vertical field of view (FOV). The current state-of-the-art sensors have 128 vertical beams and operate at 10 to 20 Hz [25]. However, low-frame-rate LiDAR sensors have existed with much higher resolutions for more than two decades. McManus et al. [27] showed that lighting-invariant visual odometry could be performed using SURF features on intensity images taken from an Autonosys 2D scanning LiDAR. This sensor produces images with a $30\degree\text{V}\times 90\degree$H FOV, with resolution $480\times 360$ pixels at 2 Hz. The frame rate is much slower, but the vertical resolution is still significantly higher (12 pixels per degree) than the OS-1 LiDAR (2.85 pixels per degree) used in this paper. This approach was extended to account for the motion distortion caused by the moving sensor [28] and used to control an offroad vehicle in closed loop [29], [30].
@@ III-A Creating Virtual Camera Views
The core idea of this pipeline is to project the two aligned point clouds into images, the format expected by SAM. The natural data format of a spinning LiDAR is equirectangular where each pixel corresponds to a constant angular offset. The natural data format of cameras is the perspective projection. Sample images created using each projection are shown in Figure 3. A larger blindspot around the base of the robot is created in the equirectangular view than in the perspective view when deviating from the path (the map view). In the following, we will use the perspective projection because the image segmentation neural networks were trained on data in this form. We define a virtual camera pose in the vehicle’s frame, $\textbf{T}_{cv}\in SE(3)$, which is used to render the images. The desired virtual camera pose is arbitrary, but aligning its position with the origin of the LiDAR sensor minimizes the amount of interpolation required in the final image. The first step is to align the point clouds in the camera frame. When a new live scan is received, an existing ICP-based localization module [31] is used to extract the relative pose of the vehicle in the map ($\textbf{T}_{vm}\in SE(3)$). The extrinsic transformation between the LiDAR sensor and the vehicle, $\textbf{T}_{vs}\in SE(3)$, is known from calibration.
@@
When a perspective projection is used, gaps in the image are created due to the distortion. These gaps are highlighted in the third row of Figure 3. Interpolating these regions is critical for the segmentation to work properly. A $3\times 3$ kernel is used to interpolate the missing pixels, but only non-zero pixels are considered. The distortion is exacerbated by any offset of the virtual camera’s origin from the LiDAR. Conceptually, the virtual camera could be colocated with the LiDAR’s position in either the teach or repeat. However, we choose to locate it in the repeat frame because the map point clouds are an accumulation of points from sequential scans. The hue-saturation-value colour model is used to colour the images by range. The $z$ coordinate in the camera frame is used for hue, mapping over a maximum range of 30 m. The saturation is the constant 255 for all pixels and the value is the LiDAR intensity of the pixel. The first column of Figure 4 shows two sample images rendered from the robot in a common frame.
@@ III-B Segmenting Regions of Interest
After rendering both images in a common frame, segmentation occurs. The pre-trained model sam_vit_b [16] is used without fine-tuning for all experiments. SAM requires a geometric prompt on the image to define the anchor point(s) of the mask. To remain as general as possible, we use a single point-prompt for each semantic mask. Prompt selection is a critical aspect of the method. Each prompt adds run time to the model so we are constrained to select a few prompts, rather than a uniform grid that covers the whole image. To select prompts, two methods were compared. The first uses the norm of the pixel difference between the live image and the map image. The top-$k$ local maxima are selected as prompts. Additionally, a minimum distance constraint is imposed to ensure greater image coverage. The second method uses the $k$-largest centroids of connected components of the 3D nearest neighbours changes projected into the camera view.
@@ III-C Detecting Changes
After independently segmenting the image pairs, the detected masks are compared. To decide if two masks capture the same object, the intersection-over-union (IoU) score is calculated between them. The bitwise AND ($\wedge$) and OR ($\vee$) operations are used to calculate the IoU of two binary masks ($b1,b2$),
@@
An IoU score of one indicates a perfect match and zero means no overlap between the individual masks. Masks with a maximum IoU of less than 0.5 are considered to be changed between the live scan and the map. A secondary 3D check is used to ensure that the point clouds intersect in 3D as well as in their projection. The remaining masks are considered unchanged. All points classified as changed in the live scan are then back-projected into 3D. The 3D bounding box, centroid, or raw points can be estimated and provided to a local planner. A corridor-constrained sampling-based planner [7] inflates the changed points to account for the robot’s footprint and avoids them.
TABLE I: Point-Wise Comparison of LaserSAM to Change Detection Baseline Algorithms.
Full Field of View Corridor Filtered
Method IoU Precision Recall IoU Precision Recall Run Time (ms)
Pixel Difference Baseline 13.7% 15.3% 55.9% 21.5% 28.0 % 48.0% 5.7 $\pm$ 1.2
LaserSAM with Pixel Difference Prompts 28.2% 39.7% 49.4 % 47.1% 68.0% 60.5% 281.3 $\pm$ 27.9
3D Difference Baseline 56.5% 95.5% 58.0% 58.6% 97.0% 59.7% 80.6 $\pm$ 13.2
LaserSAM with 3D Difference Prompts 73.3% 86.5% 82.8% 80.4% 94.4% 84.5% 356.2 $\pm$ 30.9
@@ IV Experiments
Figure 4: Change-detection masks for a sample frame. The left column shows the input images. There are three changed objects, two pedestrians and a static cone, marked in the ground truth. The segmentation results for each algorithm are shown in their respective panel.
Experiments were performed at the University of Toronto’s Institute for Aerospace Studies using a Clearpath Warthog UGV [33] equipped with an Ouster OS-1 128 beam LiDAR [25]. Data were recorded during night and day as well as during snowstorm conditions. Static obstacles placed on and around the original path and pedestrians walking near the robot were introduced as changes in the dataset. A mixture of natural and fabricated items was used to minimize the impact of the chosen materials’ reflectivity on the experiments. Additionally, duplicate items were placed near the path as part of the static scene during mapping. This demonstrates that the algorithm is performing change detection, not obstacle classification. Two different sequences were recorded with changes. The first sequence is 262 m long on relatively flat terrain. The second is 230 m through a wooded area. Changes to be detected include cones, mannequins, pedestrians and road signs.
The dataset has 12,012 live scans to use for testing. One hundred frames were randomly selected for manual annotation to provide ground truth masks. The ground-truth masks were instance segmented to allow for per-object metrics to be evaluated in addition to per-point ones. As discussed in the related work, the precise semantic meaning of a change is problem-dependent. In this dataset, only new or moved obstacles that would impact the robot’s ability to drive were marked as changed. Small scene changes such as footprints in the snow, or tracks from previous drives were marked as unchanged.
@@ IV-A Results
The pipeline was evaluated on the annotated subset of the frames. Two baselines were selected for comparison: a camera-style pixel difference method and a 3D-geometric detector with a Gaussian roughness model [6]. The pixel-wise IoU between the predicted and ground-truth segmentation masks is used for comparison. To quantify the effectiveness of each method for robot navigation, the detection performance is restricted to be within the allowable planning corridor. This planning-oriented metric is useful because errors in change detection that will not block the robot are not as critical to safe operation. The achieved results are summarized in Table I. All computational timing was performed on a laptop NVIDIA RTX A4500 GPU and an Intel i7-12800H CPU. LaserSAM was compared to both baselines, using the five most likely regions of change from each baseline as the input prompts. When prompted using pixel differences, LaserSAM significantly reduces the number of false positives, as demonstrated by the increase in precision. When prompted using 3D distances, there are fewer false positives but the mask-refining capabilities more accurately capture the whole object. This improvement is clear from the recall increase from 59.7% to 84.5% after corridor filtering.
Figure 4 shows the segmentation result for a sample frame with two moving pedestrians and a static cone. LaserSAM generates the most accurate segmentation mask, provided that it has been prompted effectively. The gaps in the traffic cone correspond to projection artifacts. While interpolation is used to generate the mask, only those pixels that correspond to a 3D measurement are retained as changes. A secondary benefit of LaserSAM is the instance groupings of points. In Figure 4, each instance is coloured according to the ground truth instance that it matches most closely. The baseline methods provide only binary labels for each point in the scan.
@@
Figure 5: A temporal sequence of semantic regions generated by SAM [16]. The shared prompt point is highlighted in the pink bulls-eye. The mask colours match between the teach and repeat.
@@
Figure 5 shows an example sequence of masks generated by SAM. When changes are detected, the map image often has a large mask corresponding to the ground or sky. The masks from SAM tend to be consistent over time which is advantageous for robot motion planning. Prompts often exist near the boundaries, which is undesirable. Once the two masks are generated, the IoU threshold filter suppresses false positives.
IV-B Closed-Loop Experiments on Warthog UGV
LaserSAM was packaged into C++ and integrated into the Visual Teach and Repeat 3 framework111github.com/utiasASRL/vtr3. The run time of LaserSAM with 3D prompting is $356\pm 30.9$ ms, which is slower than real-time for the 10 Hz OS-1 LiDAR. A performance compromise is achieved by running lidar odometry at 10 Hz but only running change detection as frequently as possible. The 2 - 3 Hz update rate for the local costmap is fast enough to avoid static and slow-moving obstacles. Much of the computation occurs on the GPU allowing odometry on the CPU to proceed in parallel. The LiDAR sensor has a near-range blindspot so the planner maintains previously detected changes in a queue, which adds temporal permanence that is not provided by LaserSAM.
Figure 6: The path of the robot in orange as it detects and avoids changes in real time using LaserSAM. The blue line represents the obstacle-free teach path.